OpenStudy (anonymous):

5(2x-7)=6-(x-3) help please?

OpenStudy (anonymous):

5(2x-7) = 6-(x-3) multiply out 5(2x-7), which is: 10x-7 = 6-(x-3) multiply out the negative sign. -x+3 so: 10x -7 = 6-x+3 10x -7 = 9-x (i just added 6+3) then just do what I did with the last question I helped.

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

You are eveything. Thank you *gets on one knee and ask you to marry me? ;)

OpenStudy (anonymous):

whoops.

OpenStudy (anonymous):

10x-35*

OpenStudy (anonymous):

10x-35=9-x*